# Data Reconstruction Attacks and Defenses: A Systematic Evaluation

### v1

@@ C.1 Local Aggregation
We first bound $\|\operatorname{\mathbb{E}}Z_{j}\|$ By Lemma C.3 we have
$\|\operatorname{\mathbb{E}}Z_{j}\|\lesssim\operatorname{\mathbb{E}}|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\|{\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I\|\lesssim\frac{BKd}{m}.$ (18)
Next we bound $\|\frac{1}{m}\sum_{j=1}^{m}Z_{j}-\operatorname{\mathbb{E}}Z_{j}\|$ with matrix Bernstein inequality. We first check the conditions of Theorem A.2:
@@
$\|Z_{j}\|\leq 2|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|(\|{\bm{w}}_{j}\|^{2}+1)\lesssim\frac{BKd\log(4Bm/\delta)}{m}$
@@
$\lesssim\operatorname{\mathbb{E}}|\sigma({\bm{w}}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}^{\top}{\bm{x}}_{i})|^{2}\|({\bm{w}}{\bm{w}}^{\top}-I)^{2}\|\lesssim\frac{B^{2}K^{2}d^{2}}{m^{2}}.$
$\mathbb{E}({\bm{a}}^{\top}Z{\bm{a}})^{2}=\mathbb{E}|\sigma({\bm{w}}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}^{\top}{\bm{x}}_{i})|^{2}({\bm{a}}^{\top}({\bm{w}}{\bm{w}}^{\top}-I){\bm{a}})^{2}\lesssim\frac{B^{2}K^{2}d^{2}}{m^{2}}.$
Then $\max_{\|{\bm{a}}\|=\|\bm{b}\|=1}(\mathbb{E}({\bm{a}}^{\top}Z\bm{b})^{2})^{1/2}\lesssim\frac{BKd}{m}$.
Moreover, $\|\operatorname{\mathbb{E}}Z_{j}\|\lesssim\frac{BKd}{m}$ by Eq. (18). Then by Theorem A.2, we have
$\|\frac{1}{m}\sum_{j=1}^{m}Z_{j}-\operatorname{\mathbb{E}}Z_{j}\|\leq\tilde{O}(\frac{BKd}{m})$
with probability $1-\frac{\delta}{2B}$. ∎
Lemma C.5.
If the learning rate $\eta=O(1)$, we have
$\begin{split}&\left\|\frac{1}{m}\sum_{j=1}^{m}\left[(p^{(1)}_{1i}-y_{1i})+(p^{(1)}_{2i}-y_{2i})\right]\left|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)(V,V,V)\right\|\\ &\quad\quad\quad\quad\quad\quad\leq\tilde{O}(\frac{B^{5/2}K}{m})\end{split}$ (19)
with probability $1-\frac{\delta}{2B}$ for any $i$.
Proof.
We define $\tilde{{\bm{Z}}}_{j}=[(p^{(1)}_{1i}-y_{1i})+(p^{(1)}_{2i}-y_{2i})]|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)(V,V,V)$ and $Z_{j}$ be the flatten of $\tilde{{\bm{Z}}}_{j}$ along the first dimension. Then we have
We first bound $\operatorname{\mathbb{E}}Z_{j}$. By Lemma C.3, we have
$\|\operatorname{\mathbb{E}}Z_{j}\|\lesssim\operatorname{\mathbb{E}}|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\|V^{\top}{\bm{w}}_{j}\|^{3}\lesssim\frac{B^{5/2}K}{m}.$ (20)
Next we bound $\|\frac{1}{m}\sum_{j=1}^{m}Z_{j}-\operatorname{\mathbb{E}}Z_{j}\|$ with matrix Bernstein inequality. We first check the conditions of Theorem A.2:
(I) We first bound the norm of $Z_{j}$. By Lemma C.3, we have
$\|Z_{j}\|\lesssim 2|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\|V^{\top}{\bm{w}}_{j}\|^{3}\leq\tilde{O}(\frac{B^{5/2}K}{m})$
with probability $1-\frac{\delta}{2Bm}$.
(II) By Lemma C.3 we have
$\lesssim\operatorname{\mathbb{E}}|\sigma({\bm{w}}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}^{\top}{\bm{x}}_{i})|^{2}\|V^{\top}{\bm{w}}\|^{6}\lesssim\frac{B^{5}K^{2}}{m^{2}}.$
(III) We have
$\max_{\|{\bm{a}}\|=\|\bm{b}\|=1}(\mathbb{E}({\bm{a}}^{\top}Z\bm{b})^{2})^{1/2}\leq\left(\mathbb{E}\left[\left\|Z\right\|^{2}\right]\right)^{1/2}\lesssim\frac{B^{5/2}K}{m}.$
Moreover, $\|\operatorname{\mathbb{E}}Z_{j}\|\lesssim\frac{B^{5/2}K}{m}$ by Eq. (20). Then by Theorem A.2, we have
$\|\frac{1}{m}\sum_{j=1}^{m}Z_{j}-\operatorname{\mathbb{E}}Z_{j}\|\leq\tilde{O}(\frac{B^{5/2}K}{m})$
with probability $1-\frac{\delta}{2B}$. ∎
Then we give lemmas showing that $p^{(1)}$ is very close to $p$ and the difference only changes a little in the reconstruction.
Lemma C.6.
If the learning rate $\eta=O(1)$, we have
$\left|z_{ki}^{(1)}-z_{ki}\right|\leq\tilde{O}(\frac{BK}{m})$ (21)
with probability $1-\frac{\delta}{BKm}$ for any $k$ and $i$.
Proof.
We have $z_{ki}=\frac{1}{m}\sum_{j=1}^{m}a_{kj}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})$ and $z_{ki}^{(1)}=\frac{1}{m}\sum_{j=1}^{m}a_{kj}^{(1)}\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})$. Note that $a_{kj}^{(1)}=a_{kj}-\frac{\eta}{m}\sum_{i=1}^{B}(p_{ki}-y_{ki})\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})$. Then by Lemma C.3 we have
$\begin{split}\left|z_{ki}^{(1)}-z_{ki}\right|&=\left|\frac{1}{m}\sum_{j=1}^{m}a_{kj}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})-\frac{1}{m}\sum_{j=1}^{m}\left(a_{kj}-\frac{\eta}{m}\sum_{i=1}^{B}(p_{ki}-y_{ki})\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right)\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})\right|\\ &\leq\frac{1}{m}\sum_{j=1}^{m}a_{kj}\left|\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})\right|+\frac{\eta}{m^{2}}\sum_{j=1}^{m}\sum_{i=1}^{B}|p_{ki}-y_{ki}|(\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i}))^{2}\\ &\quad\quad\quad\quad+\frac{\eta}{m^{2}}\sum_{j=1}^{m}\sum_{i=1}^{B}|p_{ki}-y_{ki}||\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})||\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})|\\ &\lesssim\frac{BK}{m}+\frac{B\log(2BKm^{2}/\delta)}{m}+\frac{BK\log(2BKm^{2}/\delta)}{m^{2}}\leq\tilde{O}(\frac{BK}{m})\end{split}$ (22)
with probability $1-\frac{\delta}{BKm}$. ∎
Lemma C.7.
If the learning rate $\eta=O(1)$ and $m\geq\tilde{\Omega}(BK)$, we have
$|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}|\leq\tilde{O}(\frac{B}{m})$ (23)
with probability $1-\frac{\delta}{Bm}$ for any $i$. Moreover,
$\operatorname{\mathbb{E}}|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}|^{\alpha}\leq\tilde{O}\frac{B^{\alpha}}{m^{\alpha}}$ (24)
for any $\alpha>0$ and any $i$.
Proof.
For any $i$, by Lemma C.7, $|z_{ki}^{(1)}-z_{ki}|\leq\tilde{O}(\frac{BK}{m})$ for all $k$ with probability $1-\frac{\delta}{Bm}$. Then we have
$p_{1i}^{(1)}+p_{2i}^{(1)}\leq\frac{e^{\tilde{O}(\frac{BK}{m})}\left(e^{z_{1}}+e^{z_{2}}\right)}{e^{-\tilde{O}(\frac{BK}{m})}\sum_{k=1}^{K}e^{z_{k}}}\leq e^{\tilde{O}(\frac{BK}{m})}(p_{1i}+p_{2i})$
and
$p_{1i}^{(1)}+p_{2i}^{(1)}\geq\frac{e^{-\tilde{O}(\frac{BK}{m})}\left(e^{z_{1}}+e^{z_{2}}\right)}{e^{\tilde{O}(\frac{BK}{m})}\sum_{k=1}^{K}e^{z_{k}}}\geq e^{-\tilde{O}(\frac{BK}{m})}(p_{1i}+p_{2i}).$
Thus,
$\left(e^{-\tilde{O}(\frac{BK}{m})}-1\right)\frac{2}{K}\leq p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}\leq\left(e^{\tilde{O}(\frac{BK}{m})}-1\right)\frac{2}{K}.$
Since $1-e^{-t}\leq t$ and $e^{t}-1\leq et$ for any $t\in[0,1]$, we have
$|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}|\leq\tilde{O}(\frac{B}{m})$
with probability $1-\frac{\delta}{Bm}$ for any $i$. If we choose $\delta\leq O(\frac{B^{\alpha+1}}{m^{\alpha-1}})$, we have
$\operatorname{\mathbb{E}}|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}|^{\alpha}\leq(1-\frac{\delta}{Bm})\tilde{O}(\frac{B^{\alpha}}{m^{\alpha}})+\frac{\delta}{Bm}\leq\tilde{O}(\frac{B^{\alpha}}{m^{\alpha}}).$
Lemma C.8.
If the learning rate $\eta=O(1)$ and $m\geq\tilde{\Omega}(BK)$, we have
$\left\|\frac{1}{m}\sum_{j=1}^{m}\left(p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}\right)\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|\leq\tilde{O}(\frac{Bd}{m})$ (25)
with probability $1-\frac{\delta}{2B}$ for any $i$.
Proof.
We define $Z_{j}=(p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i})\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, then we have
We first bound $\|\operatorname{\mathbb{E}}Z_{j}\|$. By Lemma C.7 we have
$\begin{split}\|\operatorname{\mathbb{E}}Z_{j}\|&\leq\operatorname{\mathbb{E}}|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}|\|\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\|\\ &\leq\left(\operatorname{\mathbb{E}}\left[|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}|^{2}\right]\right)^{1/2}\left(\operatorname{\mathbb{E}}\|\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\|^{2}\right)^{1/2}\\ &\leq\tilde{O}(\frac{Bd}{m}).\end{split}$ (26)
Next we bound $\|\frac{1}{m}\sum_{j=1}^{m}Z_{j}-\operatorname{\mathbb{E}}Z_{j}\|$ with matrix Bernstein inequality. We first check the conditions of Theorem A.2:
(I) We first bound the norm of $Z_{j}$. By Lemma C.7, we have
$\|Z_{j}\|\leq|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}||\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|(\|{\bm{w}}_{j}\|^{2}+1)\leq\tilde{O}(\frac{Bd}{m})$
with probability $1-\frac{\delta}{2Bm}$.
(II) By Lemma C.7 we have
$\leq\left(\operatorname{\mathbb{E}}\left[|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}|^{4}\right]\right)^{1/2}\left(\operatorname{\mathbb{E}}\left\|\sigma({\bm{w}}^{\top}{\bm{x}}_{i})({\bm{w}}{\bm{w}}^{\top}-I)\right\|^{4}\right)^{1/2}$
@@
$\mathbb{E}({\bm{a}}^{\top}Z{\bm{a}})^{2}$ $=\mathbb{E}|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}|^{2}({\bm{a}}^{\top}\sigma({\bm{w}}^{\top}{\bm{x}}_{i})({\bm{w}}{\bm{w}}^{\top}-I){\bm{a}})^{2}$
$\leq\left(\operatorname{\mathbb{E}}\left[|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}|^{4}\right]\right)^{1/2}\left(\operatorname{\mathbb{E}}\left\|\sigma({\bm{w}}^{\top}{\bm{x}}_{i})({\bm{w}}{\bm{w}}^{\top}-I)\right\|^{4}\right)^{1/2}$
@@
Moreover, $\|\operatorname{\mathbb{E}}Z_{j}\|\lesssim\frac{Bd}{m}$ by Eq. (26). Then by Theorem A.2, we have
@@
If the learning rate $\eta=O(1)$ and $m\geq\tilde{\Omega}(BK)$, we have
$\left\|\frac{1}{m}\sum_{j=1}^{m}\left(p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}\right)\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)(V,V,V)\right\|\leq\tilde{O}(\frac{B^{5/2}}{m})$ (27)
@@
We define $\tilde{{\bm{Z}}_{j}}=(p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)(V,V,V)$ and $Z_{j}$ be the flatten of $\tilde{{\bm{Z}}}_{j}$ along the first dimension. Then we have
@@
$\begin{split}\|\operatorname{\mathbb{E}}Z_{j}\|&\lesssim\operatorname{\mathbb{E}}|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}||\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\|V^{\top}{\bm{w}}_{j}\|^{3}\\ &\leq\left(\operatorname{\mathbb{E}}\left[|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}|^{2}\right]\right)^{1/2}\left(\operatorname{\mathbb{E}}|\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|^{2}\|V^{\top}{\bm{w}}_{j}\|^{6}\right)^{1/2}\\ &\leq\tilde{O}(\frac{B^{5/2}}{m}).\end{split}$ (28)
Next we bound $\|\frac{1}{m}\sum_{j=1}^{m}Z_{j}-\operatorname{\mathbb{E}}Z_{j}\|$ with matrix Bernstein inequality. We first check the conditions of Theorem A.2:
@@
$\|Z_{j}\|\leq|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}||\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\|V^{\top}{\bm{w}}_{j}\|^{3}\leq\tilde{O}(\frac{B^{5/2}}{m})$
@@
$\lesssim\left(\operatorname{\mathbb{E}}\left[|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}|^{4}\right]\right)^{1/2}\left(\operatorname{\mathbb{E}}\left[\sigma({\bm{w}}^{\top}{\bm{x}}_{i})\right]^{4}\left\|V^{\top}{\bm{w}}\right\|^{12}\right)^{1/2}$
@@
$\max_{\|{\bm{a}}\|=\|\bm{b}\|=1}(\mathbb{E}({\bm{a}}^{\top}Z\bm{b})^{2})^{1/2}$ $\lesssim\left(\mathbb{E}(p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i})^{2}\left[\sigma({\bm{w}}^{\top}{\bm{x}}_{i})\right]^{2}\left\|V^{\top}{\bm{w}}\right\|^{6}\right)^{1/2}$
$\leq\left(\operatorname{\mathbb{E}}\left[|p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}|^{4}\right]\right)^{1/2}\left(\operatorname{\mathbb{E}}\left[\sigma({\bm{w}}^{\top}{\bm{x}}_{i})\right]^{4}\left\|V^{\top}{\bm{w}}\right\|^{12}\right)^{1/2}$
@@
Moreover, $\|\operatorname{\mathbb{E}}Z_{j}\|\lesssim\frac{B^{5/2}}{m}$ by Eq. (28). Then by Theorem A.2, we have

### v2

@@ C.1 Local Aggregation
We first bound $\|\operatorname{\mathbb{E}}Z_{j}\|$. By Lemma C.4 and Corollary C.6 we have
$\|\operatorname{\mathbb{E}}Z_{j}\|\lesssim\operatorname{\mathbb{E}}|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\|{\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I\|\lesssim\frac{Bd}{m}.$ (17)
Next we bound $\|\frac{1}{m}\sum_{j=1}^{m}Z_{j}-\operatorname{\mathbb{E}}Z_{j}\|$ with matrix Bernstein inequality. We first check the conditions of Theorem B.2 with Lemma C.4 and Corollary C.6
(I) We first bound the norm of $Z_{j}$. BWe have
$\|Z_{j}\|\leq 2|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|(\|{\bm{w}}_{j}\|^{2}+1)\lesssim\frac{Bd\log(4Bm/\delta)}{m}$
with probability $1-\frac{\delta}{2Bm}$.
(II) We have
$\lesssim\operatorname{\mathbb{E}}|\sigma({\bm{w}}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}^{\top}{\bm{x}}_{i})|^{2}\|({\bm{w}}{\bm{w}}^{\top}-I)^{2}\|\lesssim\frac{B^{2}d^{2}}{m^{2}}.$
$\mathbb{E}({\bm{a}}^{\top}Z{\bm{a}})^{2}=\mathbb{E}|\sigma({\bm{w}}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}^{\top}{\bm{x}}_{i})|^{2}({\bm{a}}^{\top}({\bm{w}}{\bm{w}}^{\top}-I){\bm{a}})^{2}\lesssim\frac{B^{2}d^{2}}{m^{2}}.$
Moreover, $\|\operatorname{\mathbb{E}}Z_{j}\|\lesssim\frac{Bd}{m}$ by Eq. equation 17. Then by Theorem B.2, we have
with probability $1-\frac{\delta}{2B}$. ∎
Lemma C.8.
If the learning rate $\eta_{w}=O(1)$ and $\eta_{a}=O(\frac{1}{m^{2}})$, we have
$\left\|\frac{1}{m}\sum_{j=1}^{m}r_{i}^{(1)}\left|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)(V,V,V)\right\|\leq\tilde{O}(\frac{B^{5/2}}{m})$ (18)
with probability $1-\frac{\delta}{2B}$ for any $i$.
Proof.
We define $\tilde{{\bm{Z}}}_{j}=r_{i}^{(1)}|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)(V,V,V)$ and $Z_{j}$ be the flatten of $\tilde{{\bm{Z}}}_{j}$ along the first dimension. Then we have
We first bound $\operatorname{\mathbb{E}}Z_{j}$. By Lemma C.4 and Corollary C.6, we have
$\|\operatorname{\mathbb{E}}Z_{j}\|\lesssim\operatorname{\mathbb{E}}|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\|V^{\top}{\bm{w}}_{j}\|^{3}\lesssim\frac{B^{5/2}}{m}.$ (19)
Next we bound $\|\frac{1}{m}\sum_{j=1}^{m}Z_{j}-\operatorname{\mathbb{E}}Z_{j}\|$ with matrix Bernstein inequality. We first check the conditions of Theorem B.2 with Lemma C.4 and Corollary C.6:
(I) We first bound the norm of $Z_{j}$. We have
$\|Z_{j}\|\lesssim 2|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\|V^{\top}{\bm{w}}_{j}\|^{3}\leq\tilde{O}(\frac{B^{5/2}}{m})$
with probability $1-\frac{\delta}{2Bm}$.
(II) We have
$\lesssim\operatorname{\mathbb{E}}|\sigma({\bm{w}}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}^{\top}{\bm{x}}_{i})|^{2}\|V^{\top}{\bm{w}}\|^{6}\lesssim\frac{B^{5}}{m^{2}}.$
(III) We have
$\max_{\|{\bm{a}}\|=\|\bm{b}\|=1}(\mathbb{E}({\bm{a}}^{\top}Z\bm{b})^{2})^{1/2}\leq\left(\mathbb{E}\left[\left\|Z\right\|^{2}\right]\right)^{1/2}\lesssim\frac{B^{5/2}}{m}.$
Moreover, $\|\operatorname{\mathbb{E}}Z_{j}\|\lesssim\frac{B^{5/2}}{m}$ by Eq. equation 19. Then by Theorem B.2, we have
with probability $1-\frac{\delta}{2B}$. ∎
Lemma C.9.
If the learning rate $\eta_{w}=O(1)$ and $\eta_{a}=O(\frac{1}{m^{2}})$, and $m\geq\tilde{\Omega}(B)$, we have
$\left\|\frac{1}{m}\sum_{j=1}^{m}\left(r_{i}^{(1)}-r_{i}\right)\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|\leq\tilde{O}(\frac{Bd}{m})$ (20)
with probability $1-\frac{\delta}{2B}$ for any $i$.
Proof.
We define $Z_{j}=(r_{i}^{(1)}-r_{i})\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, then we have
We first bound $\|\operatorname{\mathbb{E}}Z_{j}\|$. By Lemma C.5 we have
$\begin{split}\|\operatorname{\mathbb{E}}Z_{j}\|&\leq\operatorname{\mathbb{E}}|r_{i}^{(1)}-r_{i}|\|\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\|\\ &\leq\left(\operatorname{\mathbb{E}}\left[|r_{i}^{(1)}-r_{i}|^{2}\right]\right)^{1/2}\left(\operatorname{\mathbb{E}}\|\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\|^{2}\right)^{1/2}\\ &\leq\tilde{O}(\frac{Bd}{m}).\end{split}$ (21)
Next we bound $\|\frac{1}{m}\sum_{j=1}^{m}Z_{j}-\operatorname{\mathbb{E}}Z_{j}\|$ with matrix Bernstein inequality. We first check the conditions of Theorem B.2:
@@
$\|Z_{j}\|\leq|r_{i}^{(1)}-r_{i}||\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|(\|{\bm{w}}_{j}\|^{2}+1)\leq\tilde{O}(\frac{Bd}{m})$
@@
$\leq\left(\operatorname{\mathbb{E}}\left[|r_{i}^{(1)}-r_{i}|^{4}\right]\right)^{1/2}\left(\operatorname{\mathbb{E}}\left\|\sigma({\bm{w}}^{\top}{\bm{x}}_{i})({\bm{w}}{\bm{w}}^{\top}-I)\right\|^{4}\right)^{1/2}$
@@
$\mathbb{E}({\bm{a}}^{\top}Z{\bm{a}})^{2}$ $=\mathbb{E}|r_{i}^{(1)}-r_{i}|^{2}({\bm{a}}^{\top}\sigma({\bm{w}}^{\top}{\bm{x}}_{i})({\bm{w}}{\bm{w}}^{\top}-I){\bm{a}})^{2}$
$\leq\left(\operatorname{\mathbb{E}}\left[|r_{i}^{(1)}-r_{i}|^{4}\right]\right)^{1/2}\left(\operatorname{\mathbb{E}}\left\|\sigma({\bm{w}}^{\top}{\bm{x}}_{i})({\bm{w}}{\bm{w}}^{\top}-I)\right\|^{4}\right)^{1/2}$
@@
Moreover, $\|\operatorname{\mathbb{E}}Z_{j}\|\lesssim\frac{Bd}{m}$ by Eq. equation 21. Then by Theorem B.2, we have
@@
If the learning rate $\eta_{w}=O(1)$ and $\eta_{a}=O(\frac{1}{m^{2}})$, and $m\geq\tilde{\Omega}(B)$, we have
$\left\|\frac{1}{m}\sum_{j=1}^{m}\left(r_{i}^{(1)}-r_{i}\right)\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)(V,V,V)\right\|\leq\tilde{O}(\frac{B^{5/2}}{m})$ (22)
@@
We define $\tilde{{\bm{Z}}_{j}}=(r_{i}^{(1)}-r_{i})\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)(V,V,V)$ and $Z_{j}$ be the flatten of $\tilde{{\bm{Z}}}_{j}$ along the first dimension. Then we have
@@
$\begin{split}\|\operatorname{\mathbb{E}}Z_{j}\|&\lesssim\operatorname{\mathbb{E}}|r_{i}^{(1)}-r_{i}||\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\|V^{\top}{\bm{w}}_{j}\|^{3}\\ &\leq\left(\operatorname{\mathbb{E}}\left[|r_{i}^{(1)}-r_{i}|^{2}\right]\right)^{1/2}\left(\operatorname{\mathbb{E}}|\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|^{2}\|V^{\top}{\bm{w}}_{j}\|^{6}\right)^{1/2}\\ &\leq\tilde{O}(\frac{B^{5/2}}{m}).\end{split}$ (23)
Next we bound $\|\frac{1}{m}\sum_{j=1}^{m}Z_{j}-\operatorname{\mathbb{E}}Z_{j}\|$ with matrix Bernstein inequality. We first check the conditions of Theorem B.2:
@@
$\|Z_{j}\|\leq|r_{i}^{(1)}-r_{i}||\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\|V^{\top}{\bm{w}}_{j}\|^{3}\leq\tilde{O}(\frac{B^{5/2}}{m})$
@@
$\lesssim\left(\operatorname{\mathbb{E}}\left[|r_{i}^{(1)}-r_{i}|^{4}\right]\right)^{1/2}\left(\operatorname{\mathbb{E}}\left[\sigma({\bm{w}}^{\top}{\bm{x}}_{i})\right]^{4}\left\|V^{\top}{\bm{w}}\right\|^{12}\right)^{1/2}$
@@
$\max_{\|{\bm{a}}\|=\|\bm{b}\|=1}(\mathbb{E}({\bm{a}}^{\top}Z\bm{b})^{2})^{1/2}$ $\lesssim\left(\mathbb{E}(r_{i}^{(1)}-r_{i})^{2}\left[\sigma({\bm{w}}^{\top}{\bm{x}}_{i})\right]^{2}\left\|V^{\top}{\bm{w}}\right\|^{6}\right)^{1/2}$
$\leq\left(\operatorname{\mathbb{E}}\left[|r_{i}^{(1)}-r_{i}|^{4}\right]\right)^{1/2}\left(\operatorname{\mathbb{E}}\left[\sigma({\bm{w}}^{\top}{\bm{x}}_{i})\right]^{4}\left\|V^{\top}{\bm{w}}\right\|^{12}\right)^{1/2}$
@@
Moreover, $\|\operatorname{\mathbb{E}}Z_{j}\|\lesssim\frac{B^{5/2}}{m}$ by Eq. equation 23. Then by Theorem B.2, we have